what is the future value of an annuity if the size of each payment is 2, 500, payable at the end of each quarter for five years at an interest rate of 9%


1
Expert's answer
2021-12-20T10:17:52-0500

Periodic payment = 2,500

Interest rate = 9%

Time period = 5 years

The future value of an annuity (EVA) is calculated as follows:


"FVA=P\times\frac{[(1+i)^{n} -1}{i}"

where;

P = 2500

n = 5*4 = 20

i = 9% / 4 = 2.25%


"FVA=2500\times\frac{[(1+0.025)^{20} -1}{0.025}"


"FVA=62278.8"